# A short trajectory is all you need: A transformer-based model for long-time dissipative quantum dynamics

### v1

@@ Abstract
In this communication we demonstrate that a deep artificial neural network based on a transformer model with self-attention layers can predict the long-time population dynamics of a quantum system coupled to a dissipative environment provided that the short-time population dynamics of the system is known. The transformer model developed in this work predicts the long-time dynamics of spin-boson model efficiently and very accurately across different regimes, from weak system-bath coupling to strong coupling non-Markovian regimes. Our model is more accurate than classical forecasting models, such as recurrent neural networks and is comparable to the state-of-the-art models for simulating the dynamics of quantum dissipative systems, based on kernel ridge regression. This work is the first application of the natural language processing model in the context of quantum dissipative dynamics.
@@
Simulating long-time quantum nonadiabatic dynamics of realistic complex molecular or condensed-phase systems comprised of a large number of, generally, anharmonic degrees of freedom (DOFs) remains a insurmountable challenge. In the condensed-phase the thermal reservoir or bath consists of an infinite number of degrees of freedom (DOFs) and induces an irreversible dissipation on the system altering the dynamics of the system from unitary to non-unitary.Weiss (2012); Breuer and Petruccione (2002) A plethora of numerically exact methods were developed to attack this problem including hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (2020) multi-configurational time-dependent Hartree, Wang and Thoss (2003); Meyer, Manthe, and Cederbaum (1990) quasi-adiabatic propagator path integral,Kundu and Makri (2023); Makarov and Makri (1994) time-dependent Davydov ansatz,Luo et al. (2010) time-dependent density matrix renormalization group, Ren, Shuai, and Kin-Lic Chan (2018) tensor-train split-operator Fourier transform,Greene and Batista (2017) and the stochastic equation of motion approach Yan and Shao (2016); Hsieh and Cao (2018, 2018); Han et al. (2020); Ullah et al. (2020) to name just a few. These methods, in general, are restricted to specific forms of the Hamiltonian (e.g., harmonic bath) and have a computational cost that makes long-time simulations, even with the model Hamiltonians, infeasible. Projection-based methods like the Nakajima–Zwanzig generalized quantum master equation (GQME) allow to simulate the accurate long-time dynamics at a lower computational cost, provided the memory kernel is known a priori. Brian and Sun (2021); Nakajima (1958); Zwanzig (1960); Shi and Geva (2003); Mulvihill and Geva (2021); Kelly and Markland (2013) However, calculating the numerically exact memory kernel for GQME for a general multi-level quantum system coupled to the bath is extremely difficult. Transfer tensor method (TTM) reduces the computational cost compared with GQME but it requires a set short-time system’s reduced density matrices as an input which must be generated by a numerically accurate method. Therefore, it possesses the limitations of the abovementioned methods with the exception that only short-time computationally expensive simulations are needed. Cerrillo and Cao (2014); Kananenka et al. (2016); Buser et al. (2017); Gelzinis, Rybakovas, and Valkunas (2017); Chen et al. (2020)
@@
In recent years several machine-learning (ML) based approaches for simulating long-time dynamics of quantum dissipative systems have been developed.Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021, 2022a, 2022b); Ullah et al. (2023); Ullah and Dral (2024); Lin et al. (2021); Wu et al. (2021); Akimov (2021); Zhang et al. (2024); Lin et al. (2022); Luo et al. (2009); Rodríguez et al. (2022) Such methods dramatically reduce the computational cost of quantum dynamics simulations while, in many cases, maintain high accuracy and systematic improvability achieved by increasing the number of trainable parameters and using larger data sets for model training. Several types of ML models have been explored to date including feed forward fully connected neural networks (FFNN), convolutional neural networks, (bidirectional) recurrent neural networks, convolutional recurrent neural networks, and kernel ridge regression (KRR).Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021, 2022a, 2022b); Ullah et al. (2023); Ullah and Dral (2024); Lin et al. (2021); Wu et al. (2021); Akimov (2021); Zhang et al. (2024); Lin et al. (2022); Luo et al. (2009); Rodríguez et al. (2022) For example, Rodríguez et al.Rodríguez et al. (2022) showed that KRR models outperform neural-network-based models on the relatively small data sets due to the poor scaling of KRR methods with the size of the training set.
@@
In this Communication we present a neural network model based on transformer architecture for simulating the long-time population dynamics of a two-level quantum system coupled to a infinite bosonic bath. The transformer model was introduced by Vaswani et al.Vaswani et al. (2017) in the context of natural language processing (NLP) and it soon became the state of the art for NLP. The transformer model is a neural network that learns long-range relationships or contextual information of sequential data. This is in contrast to RNNs where the range is limited by the “memory”. The transofmer architecture has become popular not only in NLP but also in computer vision and audio.Dosovitskiy et al. (2020); Radford et al. (2023) Additionally it played an important role in the development of well-known pre-trained models such as GPT (generative pre-trained transformers) and BERT (bidirectional encoder representations from transformers).Wolf et al. (2020) Recognizing their success in NLP tasks, transformers are now being recruited to address problems in computational physics. Recently the transformer architecture was used to classify light curves of astronomical objects (Astromer).Donoso-Oliva et al. (2023) In the present work we optimized and trained a transformer model to predict the long-time dynamics of a quantum dissipative system. To the best of our knowledge, this is the first application of the transformer architecture in the context of modeling the time-evolution of open quantum systems.
[…]
@@
$\hat{H}=\epsilon\hat{\sigma}_{z}+\Delta\hat{\sigma}_{x}+\hat{\sigma}_{z}\sum_{\alpha}g_{\alpha}\left(b_{\alpha}^{\dagger}+b_{\alpha}\right)+\sum_{\alpha}\omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha},$ (1)
@@
where $\hat{\sigma}_{z}=|+\rangle\langle+|-|-\rangle\langle-|$ and $\hat{\sigma}_{x}=|+\rangle\langle-|+|-\rangle\langle+|$ are the Pauli operators, $b_{\alpha}^{\dagger}\left(b_{\alpha}\right)$ is the bosonic creation (annihilation) operator of the $\alpha$th mode with the frequency $\omega_{\alpha}$, $\epsilon$ is the energetic bias, $\Delta$ is the tunneling matrix element, and $g_{\alpha}$ are the subsystem-bath coupling coefficients.
@@
where $\lambda$ is the bath reorganization energy which controls the strength of system-bath coupling and $\omega_{c}$ is the cutoff frequency which sets the primary timescale for the bath evolution $\tau_{c}=\left(\omega_{c}\right)^{-1}$. We focus on the time evolution of the expectation value of $\hat{\sigma}_{z}$ Pauli operator
@@
which is often referred to as the population difference $\left\langle\hat{\sigma}_{z}(t)\right\rangle=p_{+}(t)-p_{-}(t)$, where $p_{\pm}(t)=\operatorname{Tr}_{\mathrm{s}}$ [|±±$\left[|\pm\rangle\langle\pm|\hat{\rho}_{s}(t)\right]$. In Eq. (3) the trace is taken over the subsystem’s DOFs as denoted by ’s’ and $\hat{\rho}_{\mathrm{s}}$ is the subsystem’s reduced density operator
@@
where $\hat{\rho}(0)$ is the total system plus bath density operator and the trace is taken over the bath DOFs. The initial state of the total system is assumed to be a product state of the following form
@@
where $\hat{H}_{b}=\sum_{\alpha}\omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha}$ is the bath Hamiltonian, $Z_{\mathrm{b}}=\operatorname{Tr}_{\mathrm{b}}\left[e^{-\beta\hat{H}_{\mathrm{b}}}\right]$ is the bath partition function, $\beta=\left(k_{\mathrm{B}}T\right)^{-1}$ is the inverse temperature, and $k_{\mathrm{B}}$ is the Boltzmann constant. The initial density operator of the subsystem is chosen to be $\hat{\rho}_{\mathrm{s}}(0)=|+\rangle\langle+|$. These conditions correspond to situations where the initial preparation of the subsystem occurs quickly on the timescale of the bath relaxation.
@@
In this work the $i$th input into the ML model is given by the two vectors $\{\mathbf{x}_{i},\mathbf{t}_{i}\}$ where $\mathbf{x}_{i}=\left(x^{(1)}_{i},\ldots,x^{(T)}_{i}\right)$ is a time-ordered sequence of the expectation values of $\hat{\sigma}_{z}(t)$ Pauli operator $x^{(j)}=\langle\hat{\sigma}_{z}(t_{j})\rangle$ (or the population difference) and $\mathbf{t}_{i}$ is the vector of the corresponding times $\mathbf{t}_{i}=\left(t^{(1)}_{i},\ldots,t^{(T)}_{i}\right)$ where $T$ is the length of the input time series. Each element of the input is a pair of real-valued numbers $x^{(j)},t^{(j)}\in\mathbb{R}$. Consider a data set $\mathcal{D}=\left\{\left(\{\mathbf{x}_{i},\mathbf{t}_{i}\},\mathbf{y}_{i}\right)\right\}_{i=1}^{N}$ containing $N$ values of the series $\mathbf{x}_{i}$, the corresponding times $\mathbf{t}_{i}$ and their associated labels $\mathbf{y}_{i}$. In time-series forecasting problems, labels can describe the future states of the input sequence $\mathbf{x}_{i}$ as denoted by $\mathbf{y}_{i}=\left(x_{i}^{(T+1)},\ldots,x_{i}^{(T+m)}\right)$. In this work we train a transformer-based model to predict a single real-valued scalar quantity, the population difference of the spin-boson model $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for a single time step $m=1$ following the last input timestep. Extensions to multi-step outputs $(m>1)$ within the presented framework are straightforward.
Figure 1: Self-attention layer diagram. Each input vector $X^{(j)}\in\mathbb{R}^{d_{p}}$ of a sequence (represented by solid black circles) is projected into the (k)ey, (q)uery, and (v)alue vectors. The scaled dot-product is calculated according to Eq. (6) for each head to obtain output $Z_{i}$. There are three heads on this diagram, each is represented by a colored box. Finally, all the outputs are concatenated and projected into the output representation $Z\in\mathbb{R}^{T\times d_{p}}$ (solid purple circles).
@@
The transformer-based model employed here is inspired by the “Astromer” model,Donoso-Oliva et al. (2023) where the input corresponds to the magnitudes $\mathbf{x}$ and times $\mathbf{t}$ of a time series. The two main components of the transformer-based model are the self-attention mechanism and the positional encoding.
@@
Self attention block. Attention is one of the most crucial concepts in deep learning.Niu, Zhong, and Yu (2021) It was inspired by the processing of large data. Humans tend to focus on distinctive parts of the information rather than processing it as a whole.Corbetta and Shulman (2002) In deep learning, especially in NLP, the attention mechanisms grant the models the ability to focus on specific parts of the input that have more relevance when producing an output. In other words, each element of the output is conditioned on the selection of items in the input of the model.
@@
The classical attention mechanism in NLP is based on RNN as is called attention-based RNN. The model consists of an RNN that encodes the input and an RNN that decodes it. In between, the attention model is added which allows the attention-based RNN to focus on the parts of the input that are critical to predict the output target.Bahdanau, Cho, and Bengio (2014); Aliabadi et al. (2020) Vaswani et al. introduced the self-attention mechanism which, in contrast to the classical attention-based RNN, quantifies the importance or the relationships between parts of the input without conditioning it to the sequential order. Vaswani et al. (2017) This allows self-attention models to be trained in parallel which improves the efficiency.
@@
A self-attention layer consists of multiple self-attention heads, that compute the cosine similarities of every part of the input with itself and with any other part of the sequence. Thus, every head measures the relationship between pairs of elements of the sequence. If one element of the sequence affects the other, more attention is taken into account in the model. The self-attention layer receives as input a vector representation $X\in\mathbb{R}^{T\times d_{p}}$(see Positional Encoding below) of the input sequence and transforms the input data into $Z\in\mathbb{R}^{T\times d_{p}}$ representation that is meaningful for the model. The self-attention matrix $Z_{i}\in\mathbb{R}^{T\times d_{v}}$ for the $i$th head is a weighted sum over the input values $\mathrm{V}_{i}$ in a query-key fashion
@@
$\mathrm{Q}_{\mathrm{i}}=\mathrm{XW}_{i}^{\mathrm{q}},\quad\mathrm{K}_{\mathrm{i}}=\mathrm{XW}_{i}^{\mathrm{k}}\text{, and }\mathrm{V}_{\mathrm{i}}=\mathrm{XW}_{i}^{\mathrm{v}}\text{, }$ (7)
@@
Positional Encoding. The set of expectation values is clearly an important input information in time-series forecasting. When using the transformer model the values of corresponding times are important as well. The self-attention layer is in principle does not have any explicit temporal information about the time series provided, i.e. if $X^{(j)}$ represents a population difference in the $d_{p}$ space, different ordering of the elements of $X^{(j)}$ will produce the same attention matrix $Z$. One way to introduce the temporal information in the self-attention layer is to explicitly add it to the input $X$. This is necessary to create a representation of the time using a positional encoder (PE) and add it to the representation of the population difference in the $d_{p}$ dimensional space $X=P+PE$, with $PE\in\mathbb{R}^{T\times d_{p}}$ is the representation of the time with the positional encoding and $P\in\mathbb{R}^{T\times d_{p}}$ is the representation of the population difference, that will be generated using an FFNN. Note that the positions range from 1 to $\mathrm{T}$, where $\mathrm{T}$ is the length of the time.
The $\mathrm{PE}$ consists of trigonometric transformation based on the value of the time at different frequencies $\omega_{k}$
$PE_{j,k}=\begin{cases}\sin\left(t^{(j)}\cdot\omega_{k}\right)&k\text{ is even }\\ \cos\left(t^{(j)}\cdot\omega_{k}\right)&k\text{ is odd }\end{cases},$ (8)
@@
$\omega_{k}=\frac{1}{1000^{2k}/d_{p}}\text{. }$ (9)
Trigonometrical functions are bounded within $[-1,1]$ and capture the periodic behavior, they span wavelengths from $2\pi(k=0)$ up to $2\pi\times 1000^{2}(k=d_{p})$ as in the original implementation of the transformer model. As mentioned above this PE is non-trainable, but extension to the trainable PE exists.Moreno-Cartagena et al. (2023) Furthermore, this PE is different from the original implementation of the transformer model. In our model and in the Astromer model, the PE depends on the actual value of the time $t^{(j)}$ rather than the relative position $(j)$ in the time sequence. This choice is motivated by the Astromer architecture. Additionally, we note that in the Astromer model a factor of 1000 is used in the denominator of Eq. (9) instead of 10000 used in the original transformer model. The value of the PE at $k$th dimension is controlled by the $k$th angular frequency, this makes the smaller (larger) frequencies control the higher (smaller) dimension of the PE. In Fig. 2 we show the PE for some time inputs $\mathbf{t}$ in differing ranges of the evolution.
@@
Figure 2: Positional embedding for three different time series t. From top to bottom: $\textbf{t}=(0,0.1,0.2,\dots,4.0)$, $\textbf{t}=(8.0,8.1,8.2,\dots,12.0)$, and $\textbf{t}=(16.0,16.1,\dots,20.0)$. Colors correspond to the magnitude of the embedding given by the trigonometric functions of Eq.(8). Each time step is projected into a $d_{p}=64$ dimensional space where low dimensions have the positional information and high dimension are constant in time.

### v2

@@ Abstract
In this communication we demonstrate that a deep artificial neural network based on a transformer architecture with self-attention layers can predict the long-time population dynamics of a quantum system coupled to a dissipative environment provided that the short-time population dynamics of the system is known. The transformer neural network model developed in this work predicts the long-time dynamics of spin-boson model efficiently and very accurately across different regimes, from weak system-bath coupling to strong coupling non-Markovian regimes. Our model is more accurate than classical forecasting models, such as recurrent neural networks and is comparable to the state-of-the-art models for simulating the dynamics of quantum dissipative systems based on kernel ridge regression.
@@
Simulating long-time nonadiabatic dynamics of realistic complex molecular or condensed-phase systems comprised of a large number of, generally, anharmonic degrees of freedom (DOFs) remains a challenge. In the condensed-phase the thermal reservoir or bath induces an irreversible dissipation to the system altering its dynamics from unitary to non-unitary.Weiss (2012); Breuer and Petruccione (2002) A plethora of numerically exact methods were developed to attack this problem including hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (2020) multi-configurational time-dependent Hartree, Wang and Thoss (2003); Meyer, Manthe, and Cederbaum (1990) quasi-adiabatic propagator path integral,Kundu and Makri (2023); Makarov and Makri (1994) time-dependent Davydov ansatz,Luo et al. (2010) time-dependent density matrix renormalization group, Ren, Shuai, and Kin-Lic Chan (2018) tensor-train split-operator Fourier transform,Greene and Batista (2017) and the stochastic equation of motion approach Yan and Shao (2016); Hsieh and Cao (2018, 2018); Han et al. (2020); Ullah et al. (2020) to name just a few. These methods, in general, are restricted to specific forms of the Hamiltonian (e.g., harmonic bath) and have a computational cost that makes long-time simulations, even with the model Hamiltonians, often infeasible. Projection-based methods like the Nakajima–Zwanzig generalized quantum master equation (GQME) allow to simulate long-time quantum dynamics at a lower computational cost, provided the memory kernel is known a priori. Brian and Sun (2021); Nakajima (1958); Zwanzig (1960); Shi and Geva (2003); Mulvihill and Geva (2021); Kelly and Markland (2013) However, calculating the numerically exact memory kernel for GQME for a general multi-level quantum system coupled to a bath is extremely difficult. Transfer tensor method reduces the computational cost compared with GQME but it requires a set of short-time system’s reduced density matrices as an input which must be generated by a numerically accurate method.Cerrillo and Cao (2014); Kananenka et al. (2016); Buser et al. (2017); Gelzinis, Rybakovas, and Valkunas (2017); Chen et al. (2020) Therefore, it possesses the same limitations as the abovementioned methods with the exception that the computationally expensive simulations are needed only to generate the short-time input.
@@
In recent years several machine-learning (ML) based approaches for simulating long-time dynamics of quantum dissipative systems have been developed.Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2021, 2022a, 2022b); Ullah et al. (2023); Ullah and Dral (2024); Lin et al. (2021); Wu et al. (2021); Akimov (2021); Zhang et al. (2024); Lin et al. (2022); Luo et al. (2022); Rodríguez et al. (2022); Ullah et al. (2024) Such methods dramatically reduce the computational cost of quantum dynamics simulations while, in many cases, maintain high accuracy and systematic improvability achieved by increasing the number of trainable parameters and/or using larger data sets for model training. Several types of ML models have been explored to date including feed-forward fully-connected neural networks (FFNN),Akimov (2021); Rodríguez et al. (2022) convolutional neural networks, Herrera Rodríguez and Kananenka (2021); Ullah and Dral (2022a, b); Wu et al. (2021); Rodríguez et al. (2022) (bidirectional) recurrent neural networks (RNN), convolutional recurrent neural networks, Lin et al. (2021, 2022); Rodríguez et al. (2022) and kernel ridge regression (KRR). Ullah and Dral (2021, 2024); Zhang et al. (2024); Rodríguez et al. (2022) For example, Rodríguez et al.Rodríguez et al. (2022) showed that KRR models outperform neural-network-based (NN) models, for predicting the long-time dynamics. However, in the study of Rodríguez et al.Rodríguez et al. (2022) the data set size was restricted due to poor scaling of KRR methods with the size of the training set.
@@
In this Communication we present a neural network model, based on the transformer architecture, for simulating the long-time population dynamics of a quantum system coupled to the bath. The transformer model was originally introduced by Vaswani et al.Vaswani et al. (2017) in the context of natural language processing (NLP) and it quickly became the state of the art for NLP. The transformer neural network learns long-range relationships or contextual information from sequential data. This is in contrast to RNNs where the range is limited by the “memory”. The transformer architecture has become popular not only in NLP but also in computer vision and audio.Dosovitskiy et al. (2020); Radford et al. (2023) Additionally, recently developed pre-trained models such as GPT (generative pre-trained transformers) and BERT (bidirectional encoder representations from transformers) are based on the transformer architecture.Wolf et al. (2020) Recognizing that transformers are effectively general purpose trainable computers whose domain of applicability is not limited to NLP tasks, they are now being recruited to address various problems in computational physics. Recently the transformer architecture was used to classify light curves of astronomical objects (Astromer).Donoso-Oliva et al. (2023) In the present work we optimized and trained a transformer model to predict the long-time dynamics of a quantum dissipative system. To the best of our knowledge, this is the first application of the transformer architecture in the context of modeling the time-evolution of open quantum systems.
We choose to test our transformer implementation on a time series representing the reduced density matrix dynamics of the spin-boson model which is a cornerstone model in the study of open quantum systems due to its rich physics and widespread applicability.Leggett et al. (1987) The applications range from quantum computing,Makhlin, Schön, and Shnirman (2001) quantum phase transitions, Winter et al. (2008); Alvermann and Fehske (2008) to electron transfer in biological systems.Garg, Onuchic, and Ambegaokar (1985) The spin-boson model comprises a two-level quantum subsystem linearly coupled to a bosonic bath environment which is modeled as an ensemble of independent quantum harmonic oscillators. The total Hamiltonian in the subsystem’s basis denoted as $\{|0\rangle,|1\rangle\}$ is given by $(\hbar=1)$
@@
$\hat{H}=\epsilon\hat{\sigma}_{z}+\Delta\hat{\sigma}_{x}+\hat{\sigma}_{z}\sum_{\alpha}g_{\alpha}\left(\hat{b}_{\alpha}^{\dagger}+\hat{b}_{\alpha}\right)+\sum_{\alpha}\omega_{\alpha}\hat{b}_{\alpha}^{\dagger}\hat{b}_{\alpha},$ (1)
@@
where $\hat{\sigma}_{z}=|0\rangle\langle 0|-|1\rangle\langle 1|$ and $\hat{\sigma}_{x}=|0\rangle\langle 1|+|1\rangle\langle 0|$ are the Pauli operators, $\hat{b}_{\alpha}^{\dagger}\left(\hat{b}_{\alpha}\right)$ is the bosonic creation (annihilation) operator of the $\alpha$th mode with the frequency $\omega_{\alpha}$, $\epsilon$ is the energetic bias, $\Delta$ is the tunneling matrix element, and $g_{\alpha}$ are the subsystem-bath coupling coefficients.
@@
where $\lambda$ is the bath reorganization energy which controls the strength of system-bath coupling and $\omega_{c}$ is the cutoff frequency. We focus on the time evolution of the expectation value of $\hat{\sigma}_{z}$ Pauli operator
@@
which is often referred to as the population difference $\left\langle\hat{\sigma}_{z}(t)\right\rangle=p_{0}(t)-p_{1}(t)$, where $p_{0}(t)=\operatorname{Tr}_{\mathrm{s}}$ $\left[|0\rangle\langle 0|\hat{\rho}_{s}(t)\right]$ and $p_{1}(t)=\operatorname{Tr}_{\mathrm{s}}$ $\left[|1\rangle\langle 1|\hat{\rho}_{s}(t)\right]$. In Eq. (3) the trace is taken over the subsystem’s DOFs as denoted by ’s’ and $\hat{\rho}_{\mathrm{s}}$ is the subsystem’s reduced density operator
@@
where $\hat{\rho}(0)$ is the total system plus bath density operator and the trace is taken over the bath DOFs. The initial state of the total system is assumed to be the product state of the following form
@@
where $\hat{H}_{b}=\sum_{\alpha}\omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha}$ is the bath Hamiltonian, $Z_{\mathrm{b}}=\operatorname{Tr}_{\mathrm{b}}\left[e^{-\beta\hat{H}_{\mathrm{b}}}\right]$ is the bath partition function, $\beta=\left(k_{\mathrm{B}}T\right)^{-1}$ is the inverse temperature, and $k_{\mathrm{B}}$ is the Boltzmann constant. The initial density operator of the subsystem is chosen to be $\hat{\rho}_{\mathrm{s}}(0)=|0\rangle\langle 0|$. This initial condition corresponds to situations where the initial preparation of the subsystem occurs quickly on a timescale of the bath relaxation.
@@
In this work the $i$th input into the ML model is given by a pair of $\{\mathbf{x}_{i},\mathbf{t}_{i}\}$ where $\mathbf{x}_{i}=\left(x^{(1)}_{i},\ldots,x^{(T)}_{i}\right)$ is a time-ordered sequence of expectation values of $\hat{\sigma}_{z}(t)$ $x^{(j)}=\langle\hat{\sigma}_{z}(t_{j})\rangle$ (the population difference) and $\mathbf{t}_{i}$ is the vector of the corresponding times $\mathbf{t}_{i}=\left(t^{(1)}_{i},\ldots,t^{(T)}_{i}\right)$ where $T$ is the length of the input time series. Each element of the input vectors is a pair of real-valued numbers $x^{(j)},t^{(j)}\in\mathbb{R}$. Consider a data set $\mathcal{D}=\left\{\left(\{\mathbf{x}_{i},\mathbf{t}_{i}\},\mathbf{y}_{i}\right)\right\}_{i=1}^{N}$ containing $N$ time series $\mathbf{x}_{i}$, the corresponding times $\mathbf{t}_{i}$ and their associated labels $\mathbf{y}_{i}$. In a time-series forecasting problem, the labels can describe the future states of the input sequence $\mathbf{x}_{i}$ as denoted by $\mathbf{y}_{i}=\left(x_{i}^{(T+1)},\ldots,x_{i}^{(T+m)}\right)$. In this work we train a transformer-based NN model to predict a single real-valued scalar quantity, the population difference of the spin-boson model $\left\langle\hat{\sigma}_{z}(t)\right\rangle$ for a single time step, $m=1$, following the last input timestep, $T$. Extensions to multi-step outputs $(m>1)$ within the presented framework are straightforward.
@@
The two main components of the transformer-based model are the self-attention mechanism and the positional encoding.
@@
Self attention block. Attention is one of the most crucial concepts in deep learning.Niu, Zhong, and Yu (2021) It was inspired by the processing of large data. Humans tend to focus on distinctive parts of the information rather than processing it as a whole.Corbetta and Shulman (2002) In deep learning, especially in NLP, the attention mechanisms grant the models the ability to focus on specific parts of the input that have more relevance when producing the output. In other words, each element of the output is conditioned on the selection of items in the input of the model.
@@
The classical attention mechanism in NLP is based on RNN and is called attention-based RNN. The model consists of an RNN that encodes the input and an RNN that decodes it. In between, the attention model is added which allows the attention-based RNN to focus on the parts of the input that are critical for predicting the output target.Bahdanau, Cho, and Bengio (2014); Aliabadi et al. (2020) Vaswani et al. introduced the self-attention mechanism which, in contrast to the classical attention-based RNN, quantifies the importance of the relationships between parts of the input without conditioning it to the sequential order.Vaswani et al. (2017) This allows self-attention models to be trained in parallel which improves the efficiency.
Figure 1: Self-attention layer diagram. Each input vector $X^{(j)}\in\mathbb{R}^{d_{p}}$ of a sequence, represented by solid black circles, is projected into the (k)ey, (q)uery, and (v)alue vectors. The scaled dot product is calculated according to Eq. (6) for each head to obtain output $Z_{i}$. There are three heads on this diagram, each is represented by a colored box. Finally, all the outputs are concatenated and projected into the output representation $Z\in\mathbb{R}^{T\times d_{p}}$ (solid purple circles).
@@
A self-attention layer consists of multiple self-attention heads, that compute the cosine similarities of every part of the input with itself and with any other part of the sequence. Cosine similarity is a measure of similarity between two non-zero vectors $\mathbf{v}$ and $\mathbf{u}$ defined as $S(\mathbf{v},\mathbf{u})=\mathbf{v}\cdot\mathbf{u}||\mathbf{v}||^{-1}||\mathbf{u}||^{-1}$, where $||\mathbf{u}||$ is a norm of vector $\mathbf{u}$. Thus, every self-attention head measures the relationship between pairs of elements of the sequence. If one element of the sequence affects the other, more attention is taken into account by the model. Each self-attention head represents the data in a different form, in this way different heads can capture different correlations. The self-attention layer receives as input a vector representation $X\in\mathbb{R}^{T\times d_{p}}$(see Positional Encoding below) of the input sequence and transforms the input data into $Z\in\mathbb{R}^{T\times d_{p}}$ representation that is meaningful for the model. The self-attention matrix $Z_{i}\in\mathbb{R}^{T\times d_{v}}$ for the $i$th head is a weighted sum over the input values $\mathrm{V}_{i}$ in a query-key fashion
@@
$\mathrm{Q}_{i}=\mathrm{XW}_{i}^{\mathrm{q}},\quad\mathrm{K}_{i}=\mathrm{XW}_{i}^{\mathrm{k}}\text{, and }\mathrm{V}_{i}=\mathrm{XW}_{i}^{\mathrm{v}}\text{, }$ (7)
@@
Positional Encoding. A set of expectation values $\langle\sigma_{z}(t_{j})\rangle$ is clearly an important input information for the given problem. When using the transformer model the values of corresponding times are important as well, and is crucial when the input data is not equally spaced (which is not the case here but can be considered as a possible extension of this work). The self-attention layer in principle does not have any explicit temporal information about the time series provided, i.e. if $X^{(j)}$ represents a population difference in the $d_{p}$ space, different ordering of the elements of $X^{(j)}$ will produce the same attention matrix $Z$. One way to introduce the temporal information in the self-attention layer is to explicitly add it to the input. Then, it is necessary to create a representation of the time using a positional encoder (PE) and add it to the representation of the population difference in the $d_{p}$ dimensional space $X=P+\text{PE}$, with $\text{PE}\in\mathbb{R}^{T\times d_{p}}$ is the representation of the time with the positional encoding and $P\in\mathbb{R}^{T\times d_{p}}$ is the representation of the population difference that will be generated using an FFNN. Note that the positions range from 1 to $\mathrm{T}$, where $\mathrm{T}$ is the length of the input vector.
The $\mathrm{PE}$ consists of trigonometric transformation based on the value of the time with different frequencies $\omega_{k}$
$\text{PE}_{j,k}=\begin{cases}\sin\left(t^{(j)}\cdot\omega_{k}\right)&k\text{ is even }\\ \cos\left(t^{(j)}\cdot\omega_{k}\right)&k\text{ is odd }\end{cases},$ (8)
@@
$\omega_{k}=\frac{1}{1000^{\frac{2k}{d_{p}}}}.$ (9)
[…]
@@
Figure 2: Positional embedding for three different time series. From top to bottom: $\textbf{t}=(0,0.1,0.2,\dots,4.0)$, $\textbf{t}=(8.0,8.1,8.2,\dots,12.0)$, and $\textbf{t}=(16.0,16.1,\dots,20.0)$. Colors correspond to the magnitude of the embedding given by the trigonometric functions, Eq.(8). Each time step is projected into a $d_{p}=64$ dimensional space where low dimensions have the positional information and high dimensions are constant in time.
Data sets. The data set used here is taken from Ref. 31. We describe it here for completeness. The reduced density matrix of the spin-boson Hamiltonian is calculated for all combinations of the following parameters: $\epsilon/\Delta=\{0,1\}$,